# SHAM-OT: Rapid Subhalo Abundance Matching with Optimal Transport

### v1

@@ 8 Matching more than two functions with multi-marginal optimal transport
The Sinkhorn algorithm, extended to this setting (Benamou et al., 2015), is shown in Alg. 2. We demonstrate the multi-marginal matching between three distributions: halo mass, galaxy luminosity, and HI mass. The HI mass function is given by
@@
with $M_{*}\!=\!10^{9.94},\phi_{*}\!=\!4.5\times 10^{-3}\,\mathrm{Mpc}^{-3}\mathrm{dex}^{-1}$ and $\alpha\!=\!-1.25$ following Jones et al. (2018). Similar to before, we make sure that the lower mass limit of HI mass is chosen such that the three matched distributions are normalized consistently. The multi-marginal matching plan is shown in Fig. 6. The 1D marginals are preserved by construction.
@@ 9 Conclusions and discussion
We show that the SHAM sort-matching algorithm converges to the solution of the continuous optimal transport problem in the limit of large number of samples. By employing the discrete OT formulation, where the distributions are represented as probability mass functions instead of samples, the matching relation (transport plan) can be found with standard OT solvers at negligible compute and memory cost. We present the SHAM-OT framework, which straightforwardly extends beyond the standard SHAM hard-matching of two distributions: (i) it can naturally include scatter models via regularization, and (ii) it can match more than two distributions in the multi-marginal variant. We show the agreement of OT and SHAM in realistic demonstration examples of matching the halo mass function and the galaxy luminosity function, from both analytical prescription for the halo mass function and from a simulated halo catalogue. We demonstrate how regularized OT can be used to create scatter on the mass-luminosity relation, while satisfying the marginal constraints by construction. Finally, we show that SHAM-OT can be used to match more than two functions, through an example of matching halo mass, galaxy luminosity, and HI mass functions. The SHAM-OT provides instantaneous solutions to the abundance matching problems, which presents multiple benefits for generating simulations, especially if the matching is recomputed frequently for different parameters of luminosity or HI mass functions.
@@ Acknowledgements
We acknowledge the use of the following software packages: numpy (Van Der Walt et al., 2011), galsbi (Fischbacher et al., 2024b), POT (Flamary et al., 2021), PyCosmo (Refregier et al., 2018; Tarsitano et al., 2021; Moser et al., 2022; Hitz et al., 2024), scipy (Virtanen et al., 2020). Plots were created using matplotlib (Hunter, 2007) and trianglechain (Fischbacher et al., 2023; Kacprzak & Fluri, 2022).This project was supported in part by grant 200021_192243 from the Swiss National Science Foundation. The authors acknowledge support from SERI and the SKACH consortium.
@@
The code to reproduce this analysis will be made available upon publication under this link: https://cosmo-gitlab.phys.ethz.ch/cosmo_public/sham-ot/. The simulation data can be made available upon request.

### v2

@@ 8 Matching more than two functions with multi-marginal optimal transport
The Sinkhorn algorithm, extended to this setting (Benamou et al., 2015), is shown in Alg. 2. We demonstrate the multi-marginal matching, including scatter, between three distributions: halo mass, galaxy luminosity, and HI mass. The HI mass function is given by
@@
with $M_{*}\!=\!10^{9.94},\phi_{*}\!=\!4.5\times 10^{-3}\,\mathrm{Mpc}^{-3}\mathrm{dex}^{-1}$ and $\alpha\!=\!-1.25$ following Jones et al. (2018). Similar to before, we make sure that the lower mass limit of HI mass is chosen such that the three matched distributions are normalized consistently. The multi-marginal matching plan is shown in Fig. 6. The 1D marginals are preserved by construction, eliminating the need for complex deconvolution methods required in standard SHAM.
@@ 9 Conclusions and discussion
We show that the SHAM sort-matching algorithm converges to the solution of the continuous optimal transport problem in the limit of large number of samples. By employing the discrete OT formulation, where the distributions are represented as probability mass functions instead of samples, the matching relation (transport plan) can be found with standard OT solvers at negligible compute and memory cost. Therefore, we only need to solve a problem whose complexity depends on the number of bins in the probability mass functions – typically on the order of $O(100)$ – rather than on the number of samples, which can be as large as $O(>\!10^{6})$.
We present the SHAM-OT framework, which straightforwardly extends beyond the standard SHAM hard-matching of two distributions: (i) it can naturally include scatter models via regularization, and (ii) it can match more than two distributions in the multi-marginal variant. We show the agreement of OT and SHAM in realistic demonstration examples of matching the halo mass function and the galaxy luminosity function, from both analytical prescription for the halo mass function and from a simulated halo catalogue. We demonstrate how regularized OT can be used to create scatter on the mass-luminosity relation, while satisfying the marginal constraints by construction. Finally, we show that SHAM-OT can be used to match more than two functions, through an example of matching halo mass, galaxy luminosity, and HI mass functions. The SHAM-OT provides instantaneous solutions to the abundance matching problems, which presents multiple benefits for generating simulations, especially if the matching is recomputed frequently for different parameters of luminosity or HI mass functions.
@@ Acknowledgements
We acknowledge the use of the following software packages: numpy (Van Der Walt et al., 2011), galsbi (Fischbacher et al., 2024), POT (Flamary et al., 2021), PyCosmo (Refregier et al., 2018; Tarsitano et al., 2021; Moser et al., 2022; Hitz et al., 2025), scipy (Virtanen et al., 2020). Plots were created using matplotlib (Hunter, 2007) and trianglechain (Fischbacher et al., 2023; Kacprzak & Fluri, 2022). This project was supported in part by grant 200021_192243 from the Swiss National Science Foundation. The authors acknowledge support from SERI and the SKACH consortium.
@@
The code to reproduce this analysis is available under this link: https://cosmo-gitlab.phys.ethz.ch/cosmo_public/sham-ot/. The simulation data can be made available upon request.